OpenStudy (anonymous):

If P(A) = 0.6, P(B) = 0.55 and P(A and B) = 0.2 find the probabilities of P (A or B)=? P (not A)=? P (not B)=? p (A and (not B))=? P (not (A and B))=?

OpenStudy (anonymous):

P(AuB) = P(A)+P(B) - P(A and B)=.6 + .55-.2=.95

OpenStudy (anonymous):

P( not A)= 1-P(A)=1-.6=.4 P( not B)= 1-P(B)=1-.55=.45

OpenStudy (anonymous):

You can do the rest

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

is P(A and (not B)) = .4275?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

P(A and (not B)) = P(A) - P(A and B)=.6-.2=.4

OpenStudy (anonymous):

P (not (A and B))= 1 - P(A and B)=1-.2=.8